In the laboratory you dissolve 21.3 g of barium acetate in a volumetric flask and add water to a total volume of 375ml what is the molarity of the solution
1
Expert's answer
2020-04-03T11:49:30-0400

M (Ba(CH3COO)2) = 255.41 g/mol

n = m/M

n = 21.3/255.41 = 0.083 mol

C = n/V

C = 0.083 mol / 0.375 L

C = 0.22 M